# A Review of Privacy Metrics for Privacy-Preserving Synthetic Data Generation

### v1

@@ A Review of Privacy Metrics for Privacy-Preserving Synthetic Data Generation
Frederik Marinus Trudslev $0$-0001-5201-9326 Aalborg University fmtr@cs.aau.dk Matteo Lissandrini University of Verona matteo.lissandrini@univr.it Juan Manuel Rodriguez $0$-0002-1130-8065 Aalborg University jmro@cs.aau.dk Martin Bøgsted Aalborg University & Aalborg University Hospital martin.boegsted@rn.dk  and  Daniele Dell’Aglio Aalborg University dade@cs.aau.dk
@@ 1. Introduction
Privacy Preserving Synthetic Data Generation (PP-SDG) has emerged to produce synthetic datasets from personal data while maintaining privacy and utility. Differential privacy (DP) is the property of a PP-SDG mechanism that establishes how protected individuals are when sharing their sensitive data. It is however difficult to interpret the privacy loss ($\varepsilon$) expressed by DP. To make the actual risk associated with the privacy loss more transparent, multiple privacy metrics (PMs) have been proposed to assess the privacy risk of the data. These PMs are utilized in separate studies to assess newly introduced PP-SDG mechanisms (Guillaudeux et al., 2023; Yale et al., 2020; Yoon et al., 2020; Choi et al., 2018). Consequently, these PMs embody the same assumptions as the PP-SDG mechanism they were made to assess. Therefore, a thorough definition of how these are calculated is necessary.
@@ 2. Privacy Metrics
(2) $p:\mathcal{D}\times\mathcal{D}\rightarrow[0,1]$
@@ 2.1. Simulation-based privacy metrics
The computation of these metrics start with an adversary who takes as input the synthetic dataset $Z$ and, optionally, some background knowledge as a portion of $Y$, and it outputs a set $G$ of pairs of individuals from $Y$ and $Z$. The set $G$ denotes the guesses of the attacker, or, in other words, the outcome of the attack. This outcome is then analysed by an oracle that determines the set $C$ of correct guesses, and optionally the wrong and missing ones. Finally, the sets $C$ and $G$ are processed to compute the score.
@@ 2.2. Distance-based privacy metrics
$C_{Auth}=\{(y,y^{\prime},z)|$ $y\in Y,$
(18) $y^{\prime}=\arg\min_{y^{\prime}\in Y\setminus\{y\}}d_{E}(y,y^{\prime}),$
$z=\arg\min_{z\in Z}d_{E}(y,z)\}$
@@
(22) $\hat{y}[a]=\frac{y[a]}{w(a,y[a])+\varepsilon}\qquad\forall a\in A$
@@
$C_{ID}=\{(\hat{y},\hat{y}^{\prime},\hat{z})$ $|\hat{y}\in\hat{Y},$
$\hat{y}^{\prime}=\mathop{\arg\min}_{y^{\prime}\in\hat{Y}\setminus\{\hat{y}\}}d_{E}(\hat{y},\hat{y}^{\prime}),$
$\hat{z}=\mathop{\arg\min}_{z\in\hat{Z}}d_{E}(\hat{y},\hat{z})\}$
@@
$C_{NNDR}=\{(y^{\prime},y^{\prime\prime},z)|$ $z\in Z,$
(26) $y^{\prime}=\mathop{\arg\min}_{y^{\prime}\in Y}d_{E}(\Phi^{k}(y^{\prime}),\Phi^{k}(z))\},$
$y^{\prime\prime}=\mathop{\arg\min}_{y^{\prime\prime}\in Y\setminus\{y^{\prime}\}}d_{E}(\Phi^{k}(y^{\prime\prime}),\Phi^{k}(z))\}$
@@
$C^{\prime}_{NNDR}=\{(y^{\prime},y^{\prime\prime},z)|$ $(y^{\prime},y^{\prime\prime},z)\in C_{NNDR},$
$d_{E}(\Phi^{k}(y^{\prime}),\Phi^{k}(z))=0\}$
@@
$DCR^{\prime}(Y,Z)=1-\sigma\left(log\left(DCR(Y,Z)\right)\right),$
@@
$MDCR(Y,Z)=\frac{med\left(\{d_{E}(y,z)|y\in Y,z=\mathop{\arg\min}\limits_{z\in Z}d_{E}(y,z)\}\right)}{med\left(\{d_{E}(y,y^{\prime})|y\in Y,y^{\prime}=\mathop{\arg\min}\limits_{y^{\prime}\in Y\setminus\{y\}}d_{E}(y,y^{\prime})\}\right)},$
@@
$C^{Z}_{NNAA}=\{(y,z^{\prime},z)|$ $z\in Z,$
(29) $y=\mathop{\arg\min}_{y\in Y}d_{E}(y,z),$
$z^{\prime}=\mathop{\arg\min}_{z^{\prime}\in Z\setminus\{z\}}d_{E}(z,z^{\prime})\}$
@@
$NN$ $AA(Y,Z)=$
(30) |$\frac{|\{(y,y^{\prime},z)|(y,y^{\prime},z)\in C^{Y}_{NNAA},d_{E}(y,z)>d_{E}(y,y^{\prime})\}}{|C^{Y}_{NNAA}|}|+$
@@
$NNAA^{\prime}(Y,Z)=1-NNAA(Y,Z)$

### v2

@@ A Review of Privacy Metrics for Privacy-Preserving Synthetic Data Generation
Frederik Marinus Trudslev $0$-0001-5201-9326 Aalborg UniversityAalborgDenmark fmtr@cs.aau.dk Matteo Lissandrini $0$-0001-7922-5998 University of VeronaVeronaItaly matteo.lissandrini@univr.it Juan Manuel Rodriguez $0$-0002-1130-8065 Aalborg UniversityAalborgDenmark jmro@cs.aau.dk Martin Bøgsted $0$-0001-9192-1814 Aalborg University & Aalborg University HospitalAalborgDenmark martin.boegsted@rn.dk  and  Daniele Dell’Aglio $0$-0003-4904-2511 Aalborg UniversityAalborgDenmark dade@cs.aau.dk
Artifact Availability:
The source code is available at https://github.com/hereditary-eu/PrivEval.
@@ 1. Introduction
Privacy Preserving Synthetic Data Generation (PP-SDG) has emerged to produce synthetic datasets from personal data while maintaining privacy and utility. Differential privacy (DP) is the property of a PP-SDG mechanism that establishes how protected individuals are when sharing their sensitive data. It is however difficult to interpret the privacy budget ($\varepsilon$) expressed by DP. To make the actual risk associated with the privacy budget more transparent, multiple privacy metrics (PMs) have been proposed to assess the privacy risk of the data. These PMs are utilized in separate studies to assess newly introduced PP-SDG mechanisms (Guillaudeux et al., 2023; Yale et al., 2020; Yoon et al., 2020; Choi et al., 2018). Consequently, these PMs embody the same assumptions as the PP-SDG mechanism they were made to assess. Therefore, a thorough definition of how these are calculated is necessary.
@@ 2. Privacy Metrics
$p:\mathcal{D}\times\mathcal{D}\rightarrow[0,1]$
@@ 2.1. Simulation-based privacy metrics
The computation of these metrics start with an adversary who takes as input the synthetic dataset $Z$ and, optionally, some background knowledge as a portion of $Y$, and it captures a relation $G\subset Y\times Z$, denoting the guesses of the attacker, defined as associations between elements in $Y$ and $Z$. The relation $G$ captures the attack. This resulting pairs are then analysed by an oracle that determines the relation $C\subseteq R$ of correct guesses, and optionally the wrong and missing ones. By analysing the relations $C$ and $G$, the privacy score is computed.
@@ 2.2. Distance-based privacy metrics
$C_{Auth}=\{(y,y^{\prime},z)|y\in Y,y^{\prime}=\mathop{\arg\min}_{y^{\prime}\in Y\setminus\{y\}}d_{E}(y,y^{\prime}),z=\mathop{\arg\min}_{z\in Z}d_{E}(y,z)\}$
@@
(19) $\hat{y}[a]=\frac{y[a]}{w(a,y[a])+\varepsilon},\qquad\forall a\in A$
@@
$C_{ID}=\{(\hat{y},\hat{y}^{\prime},\hat{z})|\hat{y}\in\hat{Y},\hat{y}^{\prime}=\mathop{\arg\min}_{y^{\prime}\in\hat{Y}\setminus\{\hat{y}\}}d_{E}(\hat{y},\hat{y}^{\prime}),\hat{z}=\mathop{\arg\min}_{z\in\hat{Z}}d_{E}(\hat{y},\hat{z})\}$
@@
$C_{NNDR}=\{(y^{\prime},y^{\prime\prime},z)z\in Z,y^{\prime}=\mathop{\arg\min}_{y^{\prime}\in Y}d_{E}(\Phi^{k}(y^{\prime}),\Phi^{k}(z))\},y^{\prime\prime}=\mathop{\arg\min}_{y^{\prime\prime}\in Y\setminus\{y^{\prime}\}}d_{E}(\Phi^{k}(y^{\prime\prime}),\Phi^{k}(z))\}$
@@
$C^{\prime}_{NNDR}=\{(y^{\prime},y^{\prime\prime},z)|(y^{\prime},y^{\prime\prime},z)\in C_{NNDR},d_{E}(\Phi^{k}(y^{\prime}),\Phi^{k}(z))=0\}$
@@
(25) $DCR^{\prime}(Y,Z)=1-\sigma\left(log\left(DCR(Y,Z)\right)\right),$
@@
(26) $MDCR(Y,Z)=\frac{med\left(\{d_{E}(y,z)|y\in Y,z=\mathop{\arg\min}\limits_{z\in Z}d_{E}(y,z)\}\right)}{med\left(\{d_{E}(y,y^{\prime})|y\in Y,y^{\prime}=\mathop{\arg\min}\limits_{y^{\prime}\in Y\setminus\{y\}}d_{E}(y,y^{\prime})\}\right)},$
@@
$C^{Z}_{NNAA}=\{(y,z^{\prime},z)|z\in Z,y=\mathop{\arg\min}_{y\in Y}d_{E}(y,z),z^{\prime}=\mathop{\arg\min}_{z^{\prime}\in Z\setminus\{z\}}d_{E}(z,z^{\prime})\}$
@@
$NNAA(Y,Z)=$ |$\frac{|\{(y,y^{\prime},z)|(y,y^{\prime},z)\in C^{Y}_{NNAA},d_{E}(y,z)>d_{E}(y,y^{\prime})\}}{|C^{Y}_{NNAA}|}|+$
@@
(28) $NNAA^{\prime}(Y,Z)=1-NNAA(Y,Z)$
@@
Acknowledgements.
This work is partially supported by the HEREDITARY Project, as a part of the European Union’s Horizon Europe research and innovation programme under Grant Agreement No GA 101137074.